# Electronic stopping cross sections of tungsten to swift ions and comparisons with models

### v1

@@ Abstract
Accurate stopping power data for tungsten is crucial for ion beam analysis (IBA) techniques applied to fusion-related materials. In this work, we present new experimental measurements of the stopping power of tungsten for protons and alpha particles, addressing key gaps in fundamental databases. Our results provide a densely spaced dataset, refining the practical uncertainty limits to approximately 1.5% for protons and 4% for alpha particles. We critically compare our findings with semi-empirical and theoretical models, evaluating their performance in describing the stopping power of tungsten for light projectiles. By improving the accuracy and reliability of stopping power data, we contribute to the enhancement of the applicability of ion-beam methods for characterizing tungsten in fusion-related research. These findings contribute to the refinement of semi-empirical models and support the ongoing efforts to develop more precise theoretical frameworks for ion-solid interactions in high-Z materials.
@@ 1 INTRODUCTION
The scarcity of data in the IAEA database [6] poses challenges for evaluating existing models and limits the accuracy of IBA results in tungsten-based materials. To address this issue, the IAEA has launched a coordinated research project (CRP-F11023) aimed at, among other things, increasing the availability of high-quality experimental data for tungsten with these projectiles. This initiative seeks to enhance the applicability of ion beam analysis (IBA) in tungsten studies, contributing to the advancement of fusion reactor development.
@@ 2.1 Sample preparation and characterization
A 3-$\mu$m nominal thick foil was used in transmission experiments with the aim of stopping protons measurements. The foil is commercially available from Goodfellow with a nominal purity of 99.95%. The average areal density was experimentally determined as 6.0178(25) mg/cm2 with 0.05 at.of carbon contamination determined by measurements of resonant proton scattering. The contamination percentage was assessed by the ratio of respective signal heights, thus the result is less sensitive to energy loss models.
A sputter-deposited thin film of tungsten was used in backscattering experiments to determine the stopping to helium. The nominal thickness was 70 nm, with the experimentally determined areal atomic density at 483.63×$\times 10^{15}$ at./cm2 by proton scattering. The low energy loss of 2.0 MeV protons (below ten keV) in this film allowed for the density determination using a thin film approximation (independent of energy-loss models). No carbon nor oxygen contamination was observed in proton scattering experiments at their respective resonant energies.
@@ 2.2 Experimental procedures and data processing
Data processing for the transmission experiment followed the standards defined in a previous work [8]. The procedure consists in partially shadowing the detector’s field of view, and thus protons backscattered by a thin gold foil are recorded with a fraction of the protons going directly to the detector while the complementary fraction passes through the foil. The energy loss in the foil is then measured by the energy shift of both signals in the spectra.
@@
A Gaussian curve is adjusted for both the direct and transmitted peaks. Initially, we used a peak detection algorithm to obtain the initial value of the Gaussian center. The fine curve fit adjusts the centroid position and the width of the peaks.
@@
In backscattering experiments, the stopping information is extracted from the width of the tungsten peak in the spectra. For that, the sample was measured with a fixed detector at the 120 scattering angle under two incidence angle conditions, first at 0 and then at 60 referred to the normal vector to the surface. Thus, computing the energy losses in the way-in and way-out of the film, assuming the surface approximation in both conditions, provides us with a system equation for two stopping power values. This method is described in detail in reference [9].
@@
Systematic uncertainties were estimated by propagating the film thickness uncertainty values, while statistical uncertainties were estimated by propagating calibration and fit uncertainties together with a Monte Carlo approach using resampling of the spectra with Poisson statistics to propagate uncertainties related to counting statistics, as described in [8]. The latter is used to evaluate the robustness for peak centroid and edges determination in transmission and backscattering experiments, respectively.
@@ 2.3.1 Semi-empirical models
The semi-empirical approaches are based on the adjustment of a theory-guided function to a selection of experimental data. Due to its wide acceptance in practical applications, comparisons with new and accurate experimental data are highly desirable. One key factor is that semi-empirical approaches tend to be accurate when reliable experimental data are available. For the case of tungsten, only five experimental data sets are available in the IAEA database [6] for protons and seven are available for helium.
@@
We also compared our results with the compilation of experimental data provided by the IAEA [16]. Only a few experimental data of stopping cross sections of tungsten to protons are contained in the database, with data from the 1970s and 1980s. For the He ions, the number of experimental results in the database is higher but far from abundant. In this case, the data are from the 1960s and early 1970s. In both cases, the Uppsala group added new experimental data in 2021. Data from this group were obtained in a wide energy range, significantly contributing to the characterization of the H and He energy loss process in this material. However, additional data are needed to better assess systematic errors, and our work helps fill some gaps in the energy range covered by the Uppsala group.
@@ 2.3.2 Theoretical models
CasP (Convolution Approximation for Swift Particles) is based on a Unitary Convolution Approximation (UCA) [21], where the energy loss of a swift ion is calculated by integrating the individual contributions of electrons. This method efficiently provides stopping power predictions over a broad energy range and for diverse ion-target combinations. CasP allows replacement of UCA calculations for valence electrons by calculations based on the free-electron-gas approximations IDA (Induced Density Approximation) and TCS (Transport Cross-Section), as described in [22, 23]. Important information is that in this work calculations were performed with the option for scanning charge states of the ion, with the recombination, shell and Barkas corrections included in the model, and the number of valence electrons as 6. The stopping contribution due to projectile-electron loss (target-induced projectile ionization and excitation) and electron capture for charge exchange at charge-equilibrium conditions were also included in the model.
@@
CBETHE [19] is a recent implementation of the Bethe stopping power formula and its corrections. Includes the most recent advances in the calculations of the various corrections with extended range of applicability. It is worth mentioning that while all our data for H lie within the CBETHE validity range (E $\geq$ 750 keV), only a few He data lie within the validity range (E $\geq$ 5000 keV).
@@ 3 RESULTS AND DISCUSSION
Fig. 2 presents similar comparisons for the obtained values of tungsten stopping power for alpha particles. Once again, we observe the agreement between our data and the Uppsala results. Ours contributes with more densely spaced values, allowing for a refinement of the practical uncertainty limits to approximately 4% rms.
@@ 4 CONCLUSIONS
Our experimental results and those recently published by the Uppsala group show excellent agreement. In addition to mutual corroboration, our measurements provide a more densely spaced dataset, helping to refine the practical uncertainty limits to approximately 1.5% for protons and 4% for alpha particles within the energy range of 1000 and 5000 keV, approximately.

### v2

@@ Abstract
Accurate stopping power data for tungsten is crucial for ion beam analysis (IBA) techniques applied to fusion-related materials. In this work, we present new experimental measurements of the stopping power of tungsten for protons and alpha particles, addressing key gaps in fundamental databases. Our results provide a densely spaced dataset, refining the practical uncertainty limits to approximately 1.5% for protons and 2.0% for alpha particles. We critically compare our findings with semi-empirical and theoretical models, evaluating their performance in describing the stopping power of tungsten for light projectiles. By improving the accuracy and reliability of stopping power data, we contribute to the enhancement of the applicability of ion-beam methods for characterizing tungsten in fusion-related research. These findings contribute to the refinement of semi-empirical models and support the ongoing efforts to develop more precise theoretical frameworks for ion-solid interactions in high-Z materials.
@@ 1 INTRODUCTION
The scarcity of data in the IAEA database [6, 7] poses challenges for evaluating existing models and limits the accuracy of IBA results in tungsten-based materials. For instance, only nine experimental datasets are currently available for each of protons and helium ions, with very few data available for heavy ions. By contrast, for copper (one of the most extensively studied materials) there are 47 experimental datasets for protons, 22 for helium ions, and data for 30 different heavy ions.
To address this issue, the IAEA has launched a coordinated research project (CRP-F11023) aimed at, among other things, increasing the availability of high-quality experimental data of stopping cross-sections of tungsten to protons, helium, copper, and iodine. This initiative seeks to enhance the applicability of ion beam analysis (IBA) in tungsten studies, contributing to the advancement of fusion reactor development.
@@ 2.1 Sample preparation and characterization
A 3-$\mu$m nominal thick foil (commercially available from Goodfellow) was used in transmission experiments with the aim of stopping proton measurements. The average areal density was experimentally determined as 6.0178(25) mg/cm2 by direct mass-area ratio. The mass was obtained by a high-precision scale (Sartorius balance, model SE2) and the area was obtained using a calibrated high-resolution optical scanner. Unfortunately, our results could not take advantage of this tight precision due to foil inhomogeneities. The adopted procedure of data analysis is detailed in the next section.
The foil purity was evaluated as 95(1) at.%. with a 5 at.carbon contamination detected using resonant proton scattering. See the next section for more details on this analysis that confirmed the nominal specification.
A sputter-deposited thin film of tungsten was used in backscattering experiments to determine the stopping to helium. The nominal thickness was 70 nm, with the experimentally determined areal atomic density at 483(5)×$\times 10^{15}$ at./cm2 by proton scattering (148(2) $\mu$g/cm2 or 76.5(8) nm). The low energy loss of 2.18 MeV protons (below ten keV) in this film allowed for the density determination using a thin film approximation (independent of energy-loss models). No carbon or oxygen contamination was observed in proton scattering experiments at their respective resonant energies. No inhomogeneities were detected in sequenced measurements.
@@ 2.2 Experimental procedures and data processing
Data processing for the transmission experiment followed the standards defined in a previous work [9]. The procedure consists of partially shadowing the detector’s field of view, and thus protons backscattered by a thin gold foil are recorded with a fraction of the protons going directly to the detector while the complementary fraction passes through the foil. The energy loss in the foil is then measured by the energy shift of both signals in the spectra.
@@
A Gaussian curve is adjusted for both the direct and transmitted peaks. Firstly, we used a peak detection algorithm to obtain the initial value of the Gaussian center. Then a fine curve fit adjusts the centroid position and the width of the peaks. See Fig. 1. Due to foil roughness, the transmitted peak presents a broader and asymmetric peak. To determine the peak position, only part of the data was used in the Gaussian fit. The data interval considered in the Gaussian peak is shown by the shaded regions in Fig. 1. This procedure, inspired by other works [10, 11], is justified in terms of the determination of the most probable energy loss.
Figure 1: Typical data obtained in the transmission experiment. Shaded areas denotes the channels included in the Gaussian fit for the direct (right) and transmitted (left) peaks. Differences in the peak height is due to the restriction of assessed area in the stopping cross-section measurements.
To address the inhomogeneities in the foil, we measured proton scattering spectra at the same position where the stopping cross-section measurements were performed (alignment precision better than 0.1 mm). A broad proton beam, covering an area comparable to that used in the stopping measurements (selected as approximately 1 mm2), ensured that both measurements probed the same region of the sample, allowing for consistent comparison. Proton scattering measurements were conducted using a 3600 keV beam, and the detector placed at 170 scattering angle. The data are shown in Fg. 2.
The data analysis followed an iterative approach. Starting from the areal density determined in the previous section, we calculated the stopping cross-section based on the measured energy loss, and then used this information as input for SIMNRA [12] simulations of the scattering spectra (Fig. 2). The foil thickness was iteratively adjusted (and stopping cross-section updated accordingly) to achieve agreement between the experimental and the simulated scattering data.
This procedure converged to a foil areal density of 17.48(18)$\times$1018 atoms/cm2, corresponding to a thickness of 2.77(3) $\mu$m, which is slightly less than the nominal. The associated uncertainties were evaluated as 1% using the MultiSIMNRA method, as described in [13]. The carbon contamination in the foil was obtained consistently in this iterative process.
Figure 2: Proton scattering data used as constraint in the data processing method. This approach was used to compensate for sample inhomogeneities. The normally incident proton beam had 3600 keV kinetic energy and the detector was placed at 170 scattering angle. Carbon signal in the foil appear near the channel 500 and the small surface oxidation close to channel 600 was neglected.
@@
In backscattering experiments, the stopping information is extracted from the width of the tungsten peak in the spectra. For that, the sample was measured with a fixed detector at the 120 scattering angle under two incidence angle conditions, first at 0 and then at 60 referred to the normal vector to the surface.
Thus, computing the energy losses in the way-in and way-out of the film, assuming the surface approximation in both conditions, provides us with a system equation that reduces to:
$[\epsilon]_{in}=\frac{2\cdot\Delta E_{2}-\Delta E_{1}}{3\cdot K\cdot t}$ (1)
and
$[\epsilon]_{out}=\frac{2\cdot\Delta E_{1}-\Delta E_{2}}{3\cdot t}$ (2)
with $\Delta E_{1}=\Delta E(0^{\circ},60^{\circ})$ and $\Delta E_{2}=\Delta E(60^{\circ},0^{\circ})$. These values are attributed to the average energies in the entrance and exit path of the ion in the film, calculated here as:
$\bar{E}_{in}=E_{0}-\frac{1}{4}\Delta\bar{E}$ (3)
$\bar{E}_{out}=K\cdot E_{0}-\left(\frac{2\cdot K+1}{4}\right)\Delta\bar{E}$ (4)
with $\Delta\bar{E}=(\Delta E_{1}+\Delta E_{2})/2$. This method is described in detail in reference [14].
Figure 3: Typical data obtained in the scattering experiment. Differences in the tungsten peak width are due to different configuration of incidence and exit angles to the helium ions. The incidence energy was 2770 keV. Data obtained with incidence angle of 0 were scaled by a factor 2 so the tungsten peak presents the same height in the plot.
@@
Systematic uncertainties were estimated by propagating the film thickness and contaminant fraction uncertainties, while statistical uncertainties were estimated by propagating calibration and fit uncertainties together with a Monte Carlo approach using resampling of the spectra with Poisson statistics to propagate uncertainties related to counting statistics, as described in [9]. The latter is used to evaluate the robustness for peak centroid and edges determination in transmission and backscattering experiments, respectively.
@@ 2.3.1 Semi-empirical models
The semi-empirical approaches are based on the adjustment of a theory-guided function to a selection of experimental data. Due to its wide acceptance in practical applications, comparisons with new and accurate experimental data are highly desirable. One key factor is that semi-empirical approaches tend to be accurate when reliable experimental data are available. For the case of tungsten, only nine experimental data sets are available in the IAEA database [6] for each proton and helium.
@@
We also compared our results with the compilation of experimental data provided by the IAEA [7]. Only a few experimental data of stopping cross sections of tungsten to protons are contained in the database, with data from the 1970s and 1980s. For the He ions, the number of experimental results in the database is higher but far from abundant. In this case, the data are from the 1960s and early 1970s. In both cases, the Uppsala group added new experimental data in 2021. Data from this group were obtained in a wide energy range, significantly contributing to the characterization of the H and He energy loss processes in this material. However, additional data are needed to better assess systematic errors, and our work helps fill some gaps in the energy range covered by the Uppsala group.
@@ 2.3.2 Theoretical models
CasP (Convolution Approximation for Swift Particles) is based on a Unitary Convolution Approximation (UCA) [25], where the energy loss of a swift ion is calculated by integrating the individual contributions of electrons. This method efficiently provides stopping power predictions over a broad energy range and for diverse ion-target combinations. CasP allows the replacement of UCA calculations for valence electrons by calculations based on the free-electron-gas approximations IDA (Induced Density Approximation) and TCS (Transport Cross-Section), as described in [26, 27]. Important information is that in this work calculations were performed with the option for scanning charge states of the ion, with the recombination, shell and Barkas corrections included in the model, and the number of valence electrons as 6. The stopping contribution due to projectile-electron loss (target-induced projectile ionization and excitation) and electron capture for charge exchange at charge-equilibrium conditions was also included in the model.
@@
CBETHE [23] is a recent implementation of the Bethe stopping power formula and its corrections. It includes the most recent advances in the calculations of the various corrections with an extended range of applicability. It is worth mentioning that while all our data for H lie within the CBETHE validity range (E $\geq$ 750 keV), only a few He data lie within the validity range (E $\geq$ 5000 keV).
@@ 3 RESULTS AND DISCUSSION
The uncertainty sources considered in this work are summarized in Tab. 1, where average values of uncertainties and their respective cumulative contributions to the final uncertainty are presented.
Source of uncertainty Average Cumulative
(transmission measurements) value contribution
Energy loss measurement 0.57 0.57
(peak fit, energy and detector calibrations)
Thickness determination 1.0 1.15
Purity of the foil 1.0 1.44
SRIM data for impurities 2.0 1.5
Source of uncertainty Average Cumulative
(backscattering measurements) value contribution
Energy loss measurement 1.72 1.72
(edge fit, energy and detector calibrations)
Thickness determination 1 2.0
Purity of the film - 2.0
SRIM data for impurities - 2.0
Table 1: Uncertainty sources and their respective contributions to the final uncertainty of the stopping cross-section measurements presented in this work.
Other sources of uncertainty, such as geometrical, pulse height defect, and energy calibration, were not considered in this work since they were characterized in previous work and shown to be negligible in our setup.
@@
Fig. 5 presents similar comparisons for the obtained values of tungsten stopping power for alpha particles. Once again, we observe the agreement between our data and the Uppsala results. Ours contributes with more densely spaced values, allowing for a refinement of the practical uncertainty limits to approximately 2% rms.
@@ 4 CONCLUSIONS
Our experimental results and those recently published by the Uppsala group show excellent agreement. In addition to mutual corroboration, our measurements provide a more densely spaced dataset, helping to refine the practical uncertainty limits to approximately 1.5% for protons and 2% for alpha particles within the energy range of 1000 and 5000 keV, approximately.